OpenStudy (anonymous):

use integration by parts: integral of (4x-12)e^(-8x) dx

OpenStudy (lgbasallote):

\[\int (4x-12) e^{-8x}dx\] u = 4x - 12 du = 4dx dv = e^(-8x) v = -e^(-8x)/8 \[uv - \int vdu\] \[\implies -\frac{(4x-12)e^{-8x}}{8} - \int -\frac{4e^{-8x}}{8}dx\] does that help?

OpenStudy (lgbasallote):

@austin84 you there?

OpenStudy (anonymous):

thank you.. it did help. The book says that the final answer is (-x/2+23/16)e^(-8x)+c.do I need to repeat the integration by parts to get this answer?

OpenStudy (lgbasallote):

no..you dont need to do another integration by part...that laqst integral is a perfect integral